# Efficient Optimization of Variational Autoregressive Networks with Natural Gradient

### v1

@@ III.1 Performance on spin glasses models
In the first experiment, in order to investigate the convergence characteristics of the natural gradient method, we performed an experiment on an SK instance with $N=30$ spins, where we can still calculate the exact values of free energy through brute-force enumeration. After initializing the parameters, we employed both the natural gradient descent and Adam optimizer to minimize the variational free energy directly at $\beta=1.0$, the spin glass phase transition point. From the first row of Figure 1, it is evident that when employing the Adam optimizer, neither NADE, MADE, nor the transformer achieved convergence of the variational free energy, even after 1000 epochs. In contrast, when we switched to the natural gradient method, all three models successfully converged to the exact free energy value, particularly for NADE and MADE, which showed convergence in fewer than 100 epochs. In the second row of Figure 1, we also present the evolution of the variational free energy over running time. Although the training of the natural gradient method takes longer than Adam under the same epochs, the natural gradient can converge in a shorter overall duration, and provides an exponential speed up in computational time in this case. This experiment illustrates a substantial advantage of the natural gradient method in accelerating the convergence rate during training, as compared to the Adam optimizer.
@@
Then we followed the training procedure described in Ref. [1], implementing a temperature annealing process that starts at a high temperature. This approach facilitates the training of neural networks while effectively mitigating the potential occurrence of mode collapse phenomena. Figure 2 shows the relative error of the variational free energy for NADE, MADE and transformer, compared to the exact values. Notably, the results obtained by the natural gradient method demonstrate a superiority over those achieved by the Adam optimizer. We also observed that the performance of the transformer was inferior to that of NADE and MADE, regardless of whether the Adam optimizer or the natural gradient method was employed. We hypothesize that the variational optimization for the transformer may require more careful tuning of hyperparameters. Additionally, we provide inset plots showing the variance of the variational free energy, which equals zero when $q_{\theta}=p_{\mathrm{eq}}$, referred to as the zero variance condition. In this way, the variance serves as an indicator of the proximity between the variational distribution and the true Boltzmann distribution, providing a measure of the training quality. From the figure, it is evident that the variance associated with the natural gradient method is consistently lower than that of the Adam optimizer.
@@
Figure 3: For the random regular graph (RRG) with $N=256$ spins and degree $d=3$, we compare the relative error of the variational free energy between MADE using Adam optimizer (conventional VAN) and MADE using natural gradient (ng-VAN). The results of the mean-field ansatz is also presented (dashes line). The results are averaged over 10 independent RRG instances.
@@
Figure 5: The average computational cost per epoch under different number of parameters ($N_{p}$) and batch size ($N_{b}$) for NADE, MADE, and Transformer. The solid and dashed lines represent the results obtained using the natural gradient and the Adam optimizer respectively.
@@
We further carried out experiments on large sparse networks. For RRG with $N=256$, the free energy can still be obtained through exact tensor network contraction [39, 40, 41, 42]. Figure 3 presents the relative error of the variational free energy of MADE. In Figure 4, we also plot the evolution of the loss for one instance. The results clearly indicate that the natural gradient method exhibits a faster convergence speed compared to the Adam optimizer, while also yielding lower variational free energies across different temperature regions.
@@ III.2 Computational overhead of natural gradient
Here, we investigate the computational overhead associated with the natural gradient method. Each epoch in the training of VAN comprises three primary steps: (1) sampling from $q_{\theta}(\mathbf{s})$, (2) executing the forward pass [calculating the variational free energy], and (3) performing backpropagation for evaluating the gradient with respect to the model parameters. In the case of the natural gradient method, the backpropagation is substituted with the computation of per-sample gradients, i.e., the matrix $O$ in Eq. (15). This is a non-trivial task since deep learning frameworks usually accumulate the gradients on the entire batch of samples directly. Besides, solving Eq. (16) requires additional computational costs in double-precision floating-point format.
In Figure 5, we present the average computational cost per epoch for both the natural gradient method and the Adam optimizer under different number of parameters ($N_{p}$) and batch size ($N_{b}$). Note that for NADE, it requires $N$ forward passes to evaluate $q_{\theta}(\mathbf{s})$, while MADE and transformer only require a single forward pass. As pointed out in Ref. [36], the forward pass in NADE cannot be vectorized efficiently, leading to a situation where increases in $N_{p}$ and $N_{b}$ do not significantly affect computational time. However, with increasing $N_{p}$, the computational time for the natural gradient method increases more rapidly than that of the Adam optimizer. This difference is primarily attributed to the extra cost in the computation of per-sample gradients. In contrast, for transformer model, the computational time for backpropagation and per-sample gradient calculation are nearly identical. So the extra time required by the natural gradient method is primarily attributable to the matrix inversion process. Figure 6 depicts the average computational cost associated with solving the inverse of the FIM, which scales linearly with $N_{p}$ as expected.
@@ IV Conclusions and outlooks
In this work, we have successfully integrated an efficient computation of the natural gradient method into the VAN framework. Our approach, referred to as ng-VAN, drawing on recent developments in machine learning and VMC, utilizes a linear algebra identity to solve the inverse of the Fisher information matrix with computational cost cubic in the batch size. This enables us to train neural networks with up to $10^{6}$ parameters, which is beyond the capabilities of traditional approaches.
Our experiments on various spin glasses benchmarks suggest that compared to conventional VAN, the ng-VAN significantly improves convergence speed. Furthermore, ng-VAN also enhances the accuracy of estimating variational free energy and other relevant physical quantities. While the natural gradient method offers these advantages, it requires additional computational costs compared to first-order gradient-based algorithms such as Adam. This is primarily due to the necessity of calculating per-sample gradients as well as the inverting of the Fisher information matrix. Consequently, this leads to a trade-off between accelerating convergence speed and the increased computational time associated with each iteration. We also acknowledge the existence of other approximation methods for the Fisher information matrix, such as Kronecker-Factored Approximate Curvature (K-FAC) [43], which have been extensively investigated in VMC studies [44].
We highlight that our approach is applicable to the optimization of tensor network generative models [45, 46, 47], variational neural annealing [27], neural-enhanced Monte Carlo sampling techniques [26, 25, 48], and the recently proposed TwoBo architecture [49, 50]. Besides, ng-VAN can also be employed to investigate out-of-equilibrium systems, including chemical reaction networks [51, 52] and non-equilibrium lattice models [53].
Our PyTorch implementation is available at Ref. [54].
@@ Acknowledgements.
We thank Cheng Ye for helpful discussions and Feng Pan for kindly providing the exact free energy values of RRG. J.L. acknowledges support from Project No. 12405047 of the National Natural Science Foundation of China. Y.T. acknowledges support from Projects No. 12322501, No. 12105014 of the National Natural Science Foundation of China. P.Z. acknowledges support from Projects No. 12325501, No. 12047503, and No. 12247104 of the National Natural Science Foundation of China, and Project No. ZDRW-XX-2022-3-02 of Chinese Academy of Sciences. Part of the computation is carried out at the HPC Cluster of ITP-CAS.
Appendix A Proof of the approximation of KL divergence
Using a second-order Taylor expansion, the KL divergence can be expressed as follows, with $\theta^{\prime}=\theta+\delta\theta$:
@@ Appendix B Variational neural-network ansätze
Here, we give a brief overview of Neural Autoregressive Distribution Estimator (NADE) [35], Masked Autoencoder for Distribution Estimation (MADE) [36] and decoder-only transformer [9]. The NADE parameterizes the conditional probability as follows:
@@
where $M$ is a masking matrix that assigns negative infinity for positions corresponding to future spins, thereby enforcing the autoregressive property. Following the attention mechanism, the output is processed through a feed-forward network (FFN), which consists of two linear transformations separated by a non-linear activation function, typically ReLU:
$\mathrm{FFN}(x)=\mathrm{ReLU}(xW_{1}+b_{1})W_{2}+b_{2}$ (25)
@@
Appendix C Implementation details
In the experiment on the SK model with $N=30$, we set the hidden dimension $H=64$ for NADE. For MADE, we employ a two-layer architecture with a hidden dimension of 150. The transformer model only consists of a single transformer block, with embedding dimension $d_{k}=32$, FNN dimension $d_{ff}=128$ and number of heads $n_{\mathrm{head}}=4$. In this way, these three models exhibit a comparable number of parameters. During the annealing procedure, the training starts at $\beta=0.1$ and then subsequently decreased in increments of 0.1 until reaching $\beta=3.0$. The batch size is 1024 and the training epochs at each temperature are set to 1000. For the Adam optimizer, we choose a learning rate of $10^{-3}$. For our natural gradient method, we set a fixed learning rate of $0.1$ and a damping parameter of $\xi=10^{-3}$. In practice, we find that a fixed learning rate shows superior performance compared to the adaptive learning rate given by Eq. (11) in the main text.
In the experiment on the RRG with $N=256$, we utilize a single-layer MADE. Compared to the experiments on the SK model, the training epochs are increased to 2000, and the learning rate is set to $0.05$ for the natural gradient.

### v2

@@ III.1 Performance on variational free energy
In the first study, in order to investigate the convergence characteristics of the natural gradient method, we performed an experiment on an SK instance with $N=30$ spins, where we can still calculate the exact values of free energy through brute-force enumeration. After initializing the parameters, we employed both the natural gradient descent and Adam optimizer to minimize the variational free energy directly at $\beta=1.0$, the spin glass phase transition point. From the first row of Fig. 1, it is evident that when employing the Adam optimizer, neither NADE, MADE, nor the transformer achieved convergence of the variational free energy, even after 1000 epochs. In contrast, when we switched to the natural gradient method, all three models successfully converged to the exact free energy value, particularly for NADE and MADE, which showed convergence in fewer than 100 epochs. In the second row of Fig. 1, we also present the evolution of the variational free energy over running time. Although the training of the natural gradient method takes longer than Adam under the same epochs, the natural gradient can converge in a shorter overall duration, and provides an exponential speed up in computational time in this case. This experiment illustrates that the natural gradient method has a substantial advantage in accelerating the convergence rate during training as compared to the Adam optimizer.
@@
Then we followed the training procedure described in Ref. [1], implementing a temperature annealing process that starts at a high temperature. This approach facilitates the training of neural networks while effectively mitigating the potential occurrence of mode collapse phenomena. Figure 2 shows the relative error of the variational free energy for NADE, MADE and transformer, compared to the exact values. Notably, the results obtained by the natural gradient method demonstrate a superiority over those achieved by the Adam optimizer. We also observed that the performance of the transformer was inferior to that of NADE and MADE, regardless of whether the Adam optimizer or the natural gradient method was employed. The spectra of the FIM, as discussed in Ref. [40], shows that the FIM trace for the transformer model increases more rapidly during training compared to that of MADE and NADE. This phenomenon is associated with steeper curvature region of the loss function, suggesting that training the transformer in the VAN framework is more challenging and require more careful tuning of hyperparameters. Additionally, we provide inset plots showing the variance of the variational free energy, which equals zero when $q_{\theta}=p_{\mathrm{eq}}$ ($p_{\mathrm{eq}}=e^{-\beta E(\mathbf{s})}/Z$, assuming that mode collapse does not occur, otherwise, $Z$ may not present the original partition function of the Boltzmann distribution), referred to as the zero variance condition. In this way, the variance serves as an indicator of the proximity between the variational distribution and the true Boltzmann distribution, providing a measure of the training quality [41]. From Fig. 2, it is evident that the variance associated with the natural gradient method is consistently lower than that of the Adam optimizer.
@@
Figure 3: For the random regular graph (RRG) with $N=256$ spins and degree $d=3$, we compare the relative error of the variational free energy between MADE using Adam optimizer (conventional VAN) and MADE using natural gradient (ng-VAN). The results of the mean-field ansatz is also presented (dashed line). The results are averaged over 10 independent RRG instances.
@@
Figure 5: The relative error of the variational free energy to the exact solution of ferromagnetic Ising model on $16\times 16$ lattice with open boundary condition obtained by natural gradient (ng-VAN) and Adam optimizer (conventional VAN). The dashed line represents the critical temperature $\beta_{c}=\ln(1+\sqrt{2})/2$.
Figure 6: Top row: the average computational cost per epoch under different number of parameters ($N_{p}$) and batch size ($N_{b}$) for NADE, MADE, transformer, and PixelCNN. Here the differences in running time are attributed to the fact that natural gradient methods require the computation of per-sample gradients rather than backpropagation, along with the necessity of solving Eq. (16). Bottom row: the average computational cost of backpropagation and the computation of per-sample gradients.
@@
We further carried out experiments on large sparse networks. For RRG with $N=256$, the free energy can still be obtained through exact tensor network contraction [42, 43, 44, 45]. Figure 3 presents the relative error of the variational free energy of MADE. In Fig. 4, we also plot the evolution of the loss for one instance. The results clearly indicate that the natural gradient method exhibits a faster convergence speed compared to the Adam optimizer, while also yielding lower variational free energies across different temperature regions. It is important to note that, at first glance, the advantage of the natural gradient appear to diminish as the temperature decreases. This is primarily attributable to the annealing schedule employed during training, which helps to mitigate mode collapse. The effective performance of Adam at low temperatures depends on this annealing strategy. In contrast, if training is initiated directly at low temperatures, Adam is susceptible to being trapped in local optima and may experience mode collapse, while the natural gradient method exhibits faster convergence and remains more effective compared to Adam, see Appendix C for more results. Consequently, for ground state and combinatorial optimization problems with Ising formulations [46], the natural gradient method allows for initiation from lower temperatures and requires fewer annealing steps.
Finally, we consider the two-dimensional ferromagnetic Ising model. We employ open boundary condition without external fields, allowing for the exact solution of the free energy to be derived using the Kac-Ward formula [47]. We utilized the PixelCNN architecture [9, 10] and also follow the experimental setting outlined in Ref. [1]. In Fig. 5, we present the relative error of the variational free energy across different temperatures for system of $16\times 16$. When the training epochs are set to $1000$, the natural gradient exhibits significant advantages over the Adam optimizer. Furthermore, even with an increased training epochs of 10 0001000010\,00010 000, the natural gradient continues to outperform Adam.
@@ III.2 Computational overhead of natural gradient
Here, we investigate the computational overhead associated with the natural gradient method. Each epoch in the training of VAN comprises three primary steps: (1) sampling from $q_{\theta}(\mathbf{s})$, (2) executing the forward pass [calculating the variational free energy], and (3) performing backpropagation for evaluating the gradient with respect to the model parameters. In the case of the natural gradient method, the backpropagation is substituted with the computation of per-sample gradients, i.e., the matrix $O$ in Eq. (15). This is a nontrivial task since deep learning frameworks usually accumulate the gradients on the entire batch of samples directly. Besides, solving Eq. (16) requires additional computational costs in double-precision floating-point format.
In Fig. 6, we present the average computational cost per epoch for both the natural gradient method and the Adam optimizer under different number of parameters ($N_{p}$) and batch size ($N_{b}$). We also provide the running time for the computation of per-sample gradients and executing backpropagation. We observe that, for NADE, increases in $N_{p}$ and $N_{b}$ do not significantly affect the computational time for Adam optimizer. Furthermore, its calculation of per-sample gradients is notably more time consuming compared to other models. This is due to the fact that NADE requires $N$ forward passes to evaluate $q_{\theta}(\mathbf{s})$, similar to recurrent neural networks, while MADE, transformer, and PixelCNN necessitate only a single forward pass. Consequently, the computational inefficiency of NADE leads to lower GPU utilization as well as a more time-intensive process for computing per-sample gradients. For MADE, transformer, and PixelCNN, the time difference between backpropagation and computation of per-sample gradients is marginal, so the extra time required by the natural gradient method is primarily attributable to the matrix inversion process. Figure 7 depicts the average computational cost associated with solving the inverse of the FIM, which scales linearly with $N_{p}$ as expected.
@@ IV Conclusions and outlooks
In this work, we have successfully integrated an efficient computational approach for the natural gradient method into the VAN framework [1]. Our approach, referred to as ng-VAN, drawing on recent developments in machine learning and VMC [25, 26], utilizes a linear algebra identity to solve the inverse of the Fisher information matrix with computational cost cubic in the batch size. This enables us to train neural networks with up to $10^{6}$ parameters, which is beyond the capabilities of traditional approaches.
Our experiments on various spin glasses benchmarks suggest that compared to conventional VAN, the ng-VAN significantly improves convergence speed. Furthermore, ng-VAN also enhances the accuracy of estimating variational free energy and other relevant physical quantities. While the natural gradient method offers these advantages, it requires additional computational costs compared to first-order gradient-based algorithms such as Adam [16]. This is primarily due to the necessity of calculating per-sample gradients as well as the inverting of the Fisher information matrix. Consequently, this leads to a trade-off between accelerating convergence speed and the increased computational time associated with each iteration. We also acknowledge the existence of other approximation methods for the Fisher information matrix, such as the Kronecker-Factored Approximate Curvature [48], which have been extensively investigated in VMC studies [49].
We highlight that our approach is applicable to the optimization of tensor network generative models [50, 51, 52], variational neural annealing [53], neural-enhanced Monte Carlo sampling techniques [2, 30, 54], and the recently proposed TwoBo architecture [55, 56]. Besides, ng-VAN can also be employed to investigate out-of-equilibrium systems, including chemical reaction networks [57, 58] and nonequilibrium lattice models [59].
Our PyTorch implementation is available in Ref. [60].
@@ Acknowledgements.
We thank C. Ye for helpful discussions and F. Pan for kindly providing the exact free energy values of RRG. J.L. acknowledges support from Project No. 12405047 of the National Natural Science Foundation of China. Y.T. acknowledges support from Projects No. 12322501 and No. 12105014 of the National Natural Science Foundation of China. P.Z. acknowledges support from Projects No. 12325501, No. 12047503, and No. 12247104 of the National Natural Science Foundation of China, and Project No. ZDRW-XX-2022-3-02 of Chinese Academy of Sciences. Part of the computation is carried out at the HPC Cluster of ITP-CAS.
Appendix A Proofs and derivations in the main text
First, we give the derivation of the KL divergence approximation formula. Using a second-order Taylor expansion, the KL divergence can be expressed as follows, with $\theta^{\prime}=\theta+\delta\theta$:
@@ Appendix A Proofs and derivations in the main text
If we choose the Euclidean distance in the parameter space as the constraint: $\norm{\delta\theta}^{2}_{2}=\epsilon$, it leads to conventional gradient descent method. To illustrate this, we can reformulate the problem as
$\delta\theta^{\ast}=\operatorname*{arg\,min}_{\norm{\delta\theta}^{2}_{2}=\epsilon}F_{q}(\theta+\delta\theta).$ (21)
Note that the squared norm $\norm{\delta\theta}^{2}_{2}$ can be expressed as $\delta\theta^{\top}\delta\theta$. The corresponding Lagrangian for this optimization problem is
$\mathcal{L}(\delta\theta,\lambda)$ $=F_{q}(\theta+\delta\theta)+\lambda\left(\norm{\delta\theta}^{2}_{2}-\epsilon\right)$
$=F_{q}(\theta)+\nabla_{\theta}F_{q}(\theta)^{\top}\delta\theta+\lambda\delta\theta^{\top}\delta\theta-\lambda\epsilon.$ (22)
By setting the derivative of the Lagrangian with respect to $\delta\theta$ to zero, we obtain
$\delta\theta$ $=-\frac{1}{\lambda}\nabla_{\theta}F_{q}(\theta),$ (23)
which corresponds to the conventional gradient descent method. In this case, there is no preconditioner to rescale the gradients except for the learning rate factor $1/\lambda$, in other words, the FIM is an identity matrix.
In the main text, we utilize the linear algebra identity: $(O^{\top}O)^{-1}O^{\top}=O^{\top}(OO^{\top})^{-1}$. To establish the aforementioned identity, we begin with the fundamental matrix identity $I=(OO^{\top})(OO^{\top})^{-1}$. By multiplying both sides by $O^{\top}$, we obtain
$O^{\top}=O^{\top}(OO^{\top})(OO^{\top})^{-1}=(O^{\top}O)O^{\top}(OO^{\top})^{-1}.$ (24)
Subsequently, multiplying both sides by $(O^{\top}O)^{-1}$ yields
$(O^{\top}O)^{-1}O^{\top}=O^{\top}(OO^{\top})^{-1}.$ (25)
Similarly, we can also prove a similar result in the presence of a damping term, specifically:
$(O^{\top}O+\xi I_{N_{p}\times N_{p}})^{-1}O^{\top}=O^{\top}(OO^{\top}+\xi I_{N_{b}\times N_{b}})^{-1}.$ (26)
@@ Appendix B Variational neural-network ansätze
Here, we give a brief overview of NADE [11], MADE [12], decoder-only transformer [7], and PixelCNN [9, 10]. The NADE parameterizes the conditional probability as follows:
@@
where $M$ is a masking matrix that assigns negative infinity for positions corresponding to future spins, thereby enforcing the autoregressive property. Following the attention mechanism, the output is processed through a feed-forward network (FFN), which consists of two linear transformations separated by a nonlinear activation function, typically ReLU:
$\mathrm{FFN}(x)=\mathrm{ReLU}(xW_{1}+b_{1})W_{2}+b_{2}.$ (31)
@@
PixelCNN employs a convolutional neural network architecture to model the joint distribution of pixel intensities in an image, generating new images sequentially, one pixel at a time [9, 10]. By leveraging masked convolutions, PixelCNN ensures that the prediction for each pixel is influenced only by previously generated pixels, maintaining the autoregressive property. PixelCNN has been shown to outperform VAN that do not utilize convolutional networks on the two-dimensional lattice systems [1].
Appendix C Hyperparameters and implementation details
In our experiments, we used the default setting of Adam optimizer [16] with a learning rate of $10^{-3}$. For the natural gradient method, we found that selecting a fixed learning rate $\alpha_{\mathrm{fixed}}$ or setting $\epsilon$ in Eq. (11) within the range of $[0.01,1]$ generally yields good performance; however, the fixed learning rate tends to produce better results. If the learning rate for natural gradient optimization is excessively small, convergence can be significantly slower. Additionally, a learning rate schedulers can also be employed, such as decaying $\alpha_{\mathrm{fixed}}$ or $\epsilon$ over several epochs, to further enhance performance. Regarding the damping factor $\xi$ for regularization, we observed that setting it within a reasonable range, such as $[10^{-4},10^{-2}]$, has minimal impact on overall performance. Therefore, in our experiments, we consistently set $\xi$ to $10^{-3}$.
In the experiment on the SK model with $N=30$, we set the hidden dimension $H=64$ for NADE. For MADE, we employ a two-layer architecture with a hidden dimension of 150. The transformer model only consists of a single transformer block, with embedding dimension $d_{k}=32$, FNN dimension $d_{ff}=128$ and number of heads $n_{\mathrm{head}}=4$. In this way, these three models exhibit a comparable number of parameters. During the annealing procedure, the training starts at $\beta=0.1$ and then subsequently decreased in increments of $\delta\beta=0.1$ until reaching $\beta=3.0$. In other words, model parameters at a given $\beta$ are initialized using a well-trained model from the previous step at $\beta-\delta\beta$. The batch size is 1024 and the training epochs at each temperature are set to 1000. In the experiment on the RRG with $N=256$, we utilize a single-layer MADE. Compared to the experiments on the SK model, the training epochs are increased to 2000, and the learning rate is set to $0.05$ for the natural gradient. As for the two-dimensional Ising model, we initiate training from $\beta=0$ and gradually increase $\beta$ over $N_{\mathrm{annealing}}$ steps (training epochs) until the target temperature is achieved. We also adopted the PixelCNN architecture used in Ref. [1], which comprises 714 113714113714\,113714 113 learnable parameters. We set $N_{\mathrm{annealing}}$ to $1000$ and 10 0001000010\,00010 000, with a learning rate of $0.1$ for the natural gradient.
It is important to emphasize that the annealing schedule described above plays a critical role in the training process, particularly for the Adam optimizer. At each inverse temperature $\beta$, the model parameters are initialized using a well-trained model from the preceding step at $\beta-\delta\beta$. This annealing schedule ensures that the Adam optimizer remains effective even at higher values of $\beta$. As illustrated in Fig. 8, for an RRG with $N=256$, training directly at $\beta=1.0$ (as opposed to starting at $\beta=0.1$ and then gradually increase the inverse temperature) demonstrates that natural gradient optimization maintains robust performance. In contrast, the Adam optimizer exhibits significantly slower convergence, further highlighting the superior efficiency of natural gradient methods in regimes of low temperature.
Figure 8: Evolution of the variational free energy and the variance obtained by MADE, for the random regular graph with $N=256$ spins and $d=3$. Left: training the model directly at $\beta=1.0$. Right: training the model from $\beta=0.1$ and gradually increase the inverse temperature in increments of $\delta\beta=0.1$ to $\beta=1.0$.